Why RMS value calculations are important for parametric evaluation?


1
Expert's answer
2021-06-07T06:37:03-0400

Root Mean Square (RMS) value is the most important parameter that signifies the size.


One of the most important parameter that is used to describe the strength of an Alternating Current (AC).

RMS value of an AC voltage/current is equivalent to the DC voltage/current that produces the same heating effect when applied across an identical resistor. Hence, it is also a measure of energy content in a given signal.

In statistics, for any zero-mean random stationary signal, the RMS value is same as the standard deviation of the signal.

Current "I_{RMS}=\frac{I_m}{\sqrt2}"

Voltage"V_{RMS}=\frac{V_m}{\sqrt2}"

Vm= maximum voltage

Im=maximum current

Example : (Delay , spread ) of a multipath channel is often calculated as the RMS value of the

When two uncorrelated (or orthogonal ) signals are added together, such as noise from two independent sources, the RMS value of their sum is equal to the square-root of sum of the square of their individual RMS values.

Example velocity RMS value

"V_{rms}=\sqrt\frac{3kT}{M}"